# Time Series Anomaly Detection in the Frequency Domain with Statistical Reliability

### v1

@@ 4.2 Computing Selective $p$-values
To compute a selective $p$-value based on the conditional sampling distribution in (13), we also introduce a condition on the sufficient statistic of the nuisance parameter $\mathcal{Q}(\bm{X})$, which is defined as
$\mathcal{Q}(\bm{X})=\{\mathcal{V}(\bm{X}),\,\mathcal{U}(\bm{X})\}$ (14)
@@
This additional conditioning on $\mathcal{Q}(\bm{X})$ is a standard approach for computational tractability in the SI literature. Taking into account (13) and (14), the selective $p$-value is defined as
$p_{k}^{\text{selective}}=\mathbb{P}_{\text{H}_{0},k}(T_{k}(\bm{X})\geq T_{k}(\bm{x})\,|\,\bm{X}\in\mathcal{X}),$ (15)
where the conditioning event $\mathcal{X}$ is defined as
$\mathcal{X}=\{\bm{X}\in\mathbb{R}^{N}\,|\,\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x}),\mathcal{Q}(\bm{X})=\mathcal{Q}(\bm{x})\}.$ (16)
The subspace $\mathcal{X}$ is restricted to one-dimensional data space in $\mathbb{R}^{N}$ (Lee et al., 2016; Liu et al., 2018), as stated in the following theorem.
Theorem 1
The set $\mathcal{X}$ in (16) can be rewritten using a scalar parameter $z\in\mathbb{R}$ as
$\mathcal{X}=\{\bm{X}=\bm{a}+\bm{b}z\,|\,z\in\mathcal{Z}\},$ (17)
where $\bm{a},\bm{b}\in\mathbb{R}^{N}$, and the truncation region $\mathcal{Z}$ are defined as
$\bm{a}=\mathcal{U}(\bm{x}),~{}~{}~{}\bm{b}=\mathcal{V}(\bm{x}),$ (18)
$\mathcal{Z}=\{z\in\mathbb{R}\,|\,\mathcal{A}(\bm{a}+\bm{b}z)=\mathcal{A}(\bm{x})\}.$ (19)
The proof of Theorem 1 is deferred to Appendix B.1. Let us denote a random variable $Z=T_{k}(\bm{X})\in\mathbb{R}$ and its observation $z^{\text{obs}}=T_{k}(\bm{x})\in\mathbb{R}$ that satisfy $\bm{X}=\bm{a}+\bm{b}Z$ and $\bm{x}=\bm{a}+\bm{b}z^{\text{obs}}$, respectively. The selective $p$-value in (15) can be rewritten as
$p_{k}^{\text{selective}}=\mathbb{P}_{\text{H}_{0},k}(Z\geq z^{\text{obs}}\,|\,Z\in\mathcal{Z}).$ (20)
Since the uncnoditional variable $Z\sim\chi(\mathrm{tr}(P_{k}))$ under the null hypothesis, the conditional variable $Z\,|\,Z\in\mathcal{Z}$ follows a truncated $\chi$-distribution with $\mathrm{tr}(P_{k})$ degrees of freedom and truncation region $\mathcal{Z}$. Once the truncation region $\mathcal{Z}$ is identified, the selective $p$-value in (20) can be computed as
$p_{k}^{\text{selective}}=1-{F_{\text{tr}\left(P_{k}\right)}^{\text{cdf}}}^{\mathcal{Z}}\,(z^{\text{obs}}),$
@@
where ${F_{\text{tr}(P_{k})}^{\text{cdf}}}^{\mathcal{Z}}$ is the cumulative distribution function of the truncated $\chi$-distribution with $\mathrm{tr}(P_{k})$ degrees of freedom and the truncation region $\mathcal{Z}$. A schematic illustration of the SI framework is shown in Figure 5.
Figure 5: Schematic illustration of the SI framework. A point in the data space $\mathbb{R}^{N}$ corresponds to a sequence with length $N$. The darkly shaded regions in the data space indicate that, if we input a point in these regions into the algorithm $\mathcal{A}$, the CP candidates are the same as $\bm{\mathcal{T}}^{\rm det}$ obtained from the observed sequence $\bm{x}$. By conditioning on these regions and $\mathcal{Q}(\bm{X})$, the conditional sampling distribution of the test statistic $T_{k}(\bm{X})$ is represented as a truncated $\chi$-distribution. Selective $p$-values are defined based on the tail probability of such a truncated $\chi$-distribution.
Theorem 2
The selective $p$-value satisfies the property of a valid $p$-value:
@@
$\mathbb{P}_{\rm{H}_{0},k}(p_{k}^{\rm{selective}}\leq\alpha)=\alpha,\,\forall\alpha\in[0,1].$
The proof of Theorem 2 is deferred to Appendix B.2. This theorem guarantees that the type I error rate can be controlled at any significance level $\alpha$ by using the selective $p$-value.
@@
Several methods for computing selective $p$-values have been proposed in the SI community. In this study, we adopt the method based on parametric programming (Duy & Takeuchi, 2022) for the identification of the truncation region $\mathcal{Z}$. For further details, refer to Appendix C.
@@ 5.2 Synthetic Data Experiments
Experimental setup. In all synthetic experiments, we set window size $M\in\{512,1024\}$, the number of frequencies $D=\left\lfloor\frac{M}{2}\right\rfloor+1$, the length of sequence $N=M\cdot T$, where $T$ was specified for each experiment, the sampling rate $f_{s}=20480$, and each element of mean vector $\bm{s}$ as
@@
where frequencies $d_{1},d_{2},d_{3}\in\{0,...,D-1\}$ were randomly selected without replacement for each simulation, $A_{n}^{(d)}$ was defined for each experiment, and $\omega^{(d)}\in\left\{2\pi(\frac{f_{s}}{M})d\,\big{|}\,d=0,\dots,\left\lfloor\frac{M}{2}\right\rfloor\right\}$. We used BIC for the choice of penalty parameters $\bm{\beta}$ and $\gamma$ as indicated in Appendix A, and set the parameters of simulated annealing as $c_{0}^{+}=1000,\lambda^{+}=1.5,\chi(c_{0})=0.5$ and $\lambda=0.8$ in Section 3.4. After detecting CP candidates, a CP location $\tau_{k}^{\text{det}}$ randomly selected from $\bm{\tau}^{\text{det}}$ was tested at the significance level $\alpha=0.05$.
In the experiments conducted to evaluate the control of type I error rate, we generated $1000$ null sequences, which did not contain true CPs in the frequency domain, $\bm{x}=(x_{1},\dots,x_{N})^{\top}\sim\mathcal{N}\left(\bm{s},\sigma^{2}I_{N}\right)$, where $A_{n}^{(d)}=A^{(d)}$ was randomly sampled from $[0,1)$ for $d$ in each simulation, and $\sigma=1$, for each $T\in\{40,60,80,100\}$.
@@
$A_{n}^{(d)}=\begin{cases}A^{(d)}&\left(1\leq t\leq M\cdot t_{1}^{(d)}\right)\\ \vspace{-3mm}\\ A^{(d)}+\Delta&\left(M\cdot t_{1}^{(d)}+1\leq t\leq M\cdot t_{2}^{(d)}\right)\\ \vspace{-3mm}\\ A^{(d)}+2\Delta&\left(M\cdot t_{2}^{(d)}+1\leq t\leq T\right)\end{cases},$
with $A^{(d_{1})},A^{(d_{2})},A^{(d_{3})}\in[0,1)$ which were randomly sampled in each simulation, $\left(t_{1}^{(d_{1})},t_{1}^{(d_{2})},t_{1}^{(d_{3})}\right)=(18,20,22),\left(t_{2}^{(d_{1})},t_{2}^{(d_{2})},t_{2}^{(d_{3})}\right)=(38,40,42)$, an intensity of the change $\Delta\in\{0.04,0.08,0.12,0.16\}$ and $\sigma=1$, for $T=60$. In each case, we ran $1000$ trials. Since we tested only when a CP candidate location was correctly detected, the power was defined as follows
@@
We considered the CP candidate location $\tau_{k}^{\text{det}}$ to be correctly detected if it satisfied the following two conditions:
  • The set $\mathcal{D}^{\text{det}}$ of frequencies containing at least one CP candidate was a subset of $\{d_{1},d_{2},d_{3}\}$.
  • For $\mathcal{D}^{\text{det}}$ satisfying the above condition, either $\underset{d\in\mathcal{D}^{\text{det}}}{\min}{t_{1}^{(d)}}\leq\tau_{k}^{\text{det}}\leq\underset{d\in\mathcal{D}^{\text{det}}}{\max}{t_{1}^{(d)}}$ or $\underset{d\in\mathcal{D}^{\text{det}}}{\min}{t_{2}^{(d)}}\leq\tau_{k}^{\text{det}}\leq\underset{d\in\mathcal{D}^{\text{det}}}{\max}{t_{2}^{(d)}}$ held, that is, $\tau_{k}^{\text{det}}$ was detected within the true CP locations for the frequencies in $\mathcal{D}^{\text{det}}$.
Experimental results. The results of experiments regarding the control of the type I error rate are shown in Figure 6. The Proposed, OC, and Bonferroni successfully controlled the type I error rate below the significance level, whereas the OptSeg-SI, OptSeg-SI-oc, and Naive could not. That was because the OptSeg-SI and OptSeg-SI-oc used $p$-values conditioned only on the dynamic programming algorithm, excluding the conditioning on simulated annealing, and the Naive employed conventional $p$-values without conditioning. Since the OptSeg-SI, OptSeg-SI-oc, and Naive failed to control the type I error rate, we omitted the analysis of their power. The results of power experiments are shown in Figure 7. Based on these results, the Proposed was the most powerful of all methods that controlled the type I error rate. The power of the OC was lower than that of the Proposed due to redundant conditions (see Appendix C for details). Furthermore, the Bonferroni method had the lowest power because it was a highly conservative approach that accounted for the huge number of all possible hypotheses. Additionally, we provide the computational time of the Proposed in both experiments and the information on the computer resources in Appendix D.2.
(a) $M=512$
(b) $M=1024$
Figure 6: Type I Error Rate
(a) $M=512$
(b) $M=1024$
Figure 7: Power
Robustness of type I error rate control. We also conducted the following experiments to investigate the robustness of the Proposed in terms of type I error rate control.
@@
These details and results are shown in Appendix D.3.
@@ 5.3 Real Data Experiments
To demonstrate the practical applicability of the Proposed, we applied the Proposed, OC, and Naive to a real-world dataset. We used the set No.2 of the IMS bearing dataset, which is provided by the Center for Intelligent Maintenance Systems (IMS), University of Cincinnati (Qiu et al., 2006) and is available from the Prognostic data repository of NASA (Lee et al., 2007). The experimental apparatus consisted of four identical bearings installed on a common shaft, driven at a constant rotation speed by an AC motor under applied radial loading. In this dataset, the vibration signals were measured using accelerometers until the outer race of bearing 1 failed at the end of the experiment, as shown in Figure 8. The analysis for the time signal of bearing 1 in the frequency domain had revealed that anomalies were detected in the harmonics of the characteristic frequency ($236$ Hz) associated with the outer race failure (Ball Pass Frequency Outer race, BPFO) on 3–4 days (Gousseau et al., 2016). Based on this previous study, we conducted CP candidate selection for the sensor data of bearing 1 in the frequency domain (1400–4000 Hz) for two periods: 0.25–2.25 days when no anomalies existed and 2.25–4.25 days when the BPFO harmonics exhibited anomalous patterns. Subsequently, we tested the detected CP candidate locations to evaluate whether each of them was a genuine anomaly. Since the signal with 20480 samples per second had been recorded every 10 minutes, we computed the DFT of $M=1024$ consecutive points in the 20480 samples and repeated the procedure $T=288$ times. Additionally, the variance $\sigma^{2}$ was estimated from the data on 0–0.25 days that was not used in all experiments. The results for the signal of bearing 1 on 0.25–2.25 days and 2.25–4.25 days are shown in Figure 9. In panel (a), the time variation of a frequency spectrum (1920 Hz) where a CP candidate location was falsely detected is shown for the period of 0.25–2.25 days when the frequency anomaly did not actually exist. It shows that $p$-values of the Proposed and OC are above the significance level $0.05$, and therefore the result provides the validity of the inference, while $p$-value of the Naive is too small. In panel (b), the time variations of the 8th and 15th harmonics of the BPFO where CP candidate locations were correctly detected are presented for the period of 2.25–4.25 days when the frequency anomaly truly existed. In this case, $p$-values of the Proposed are below the significance level $\frac{0.05}{2}=0.025$ decided by Bonferroni correction, thus it indicates that the inference is valid. In contrast, $p$-values of the OC are too large, due to the loss of power caused by the redundant conditions. In addition, since even the time sequences of the healthy bearings 2, 3, and 4 had been reported to indicate frequency characteristics associated with the outer race fault in bearing 1 (Gousseau et al., 2016), we performed the same analysis for the three signals. The results are shown in Appendix D.4.
(a) Bearing 1
(b) Bearing 2
(c) Bearing 3
(d) Bearing 4
Figure 8: Vibration signals of the four bearings.
@@
(a) Inference on a falsely detected CP candidate location for 1920 Hz (around the 8th harmonic) on 0.25–2.25 days
@@
(b) Inferences on truely detected CP candidate locations for 1880 Hz (the 8th harmonic, above) and 3540 Hz (the 15th harmonic, below) on 2.25–4.25 days
@@
6 Conclusion
In this paper, we developed a statistical inference method to quantify the reliability of detected CP locations in the frequency domain. Our proposed framework contributes to various fields where time-frequency analysis is widely employed, such as condition monitoring of machine systems using vibration, electrical, and acoustic signals, and medical diagnosis based on biosignals. We conducted comprehensive experiments on both synthetic and real-world datasets. The results theoretically confirmed that our method provided an unbiased evaluation based on SI framework and demonstrated its superior performance compared to existing mothods. As future works, we will extend our method to the case of multi-dimensional sequences. For instance, analyzing different types of time series obtained from multiple sensors would reveal anomalies unattainable through the univariate approach, and reliability guarantee for the detections also provide a valuable future contribution.

### v2

@@ 4.2 Computing Selective $p$-value
To compute a selective $p$-value based on the conditional sampling distribution in (15), we introduce an additional condition on the sufficient statistic of the nuisance parameter $\mathcal{Q}(\bm{X})$, which is defined as
$\mathcal{Q}(\bm{X})=\left(\mathcal{V}(\bm{X}),\,\mathcal{U}(\bm{X})\right)\in\mathbb{R}^{N}\times\mathbb{R}^{N},$ (16)
@@
This additional conditioning on $\mathcal{Q}(\bm{X})$ is a standard approach for computational tractability in the SI literature (Loftus & Taylor, 2015). Based on the additional conditioning on $\mathcal{Q}(\bm{X})$, the following theorem tells that the conditional distribution of the test statistic can be represented as a truncated $\chi$-distribution.
Theorem 1.
Consider a random sequence $\bm{X}\sim\mathcal{N}(\bm{s},\sigma^{2}I_{N})$ and an observed sequence $\bm{x}$. Let $\mathcal{A}({\bm{X}})$ and $\mathcal{A}({\bm{x}})$ be the detected CP candidates, by applying a CP candidate selection algorithm to $\bm{X}$ and $\bm{x}$, respectively. Let $P_{k}\in\mathbb{R}^{N\times N}$ be a projection matrix depending on $\mathcal{A}({\bm{x}})$, and consider a test statistic in the form of $T_{k}(\bm{X})=\sigma^{-1}||P_{k}\bm{X}||$. Furthermore, define the nuisance parameter $\mathcal{Q}(\bm{X})=(\mathcal{V}(\bm{X}),\,\mathcal{U}(\bm{X}))$ as in (16).
Then, under the null hypothesis that the norm of the true signal vector $\bm{s}$ projected by $P_{k}$ is zero (i.e., $\|P_{k}\bm{s}\|=0$), the conditional distribution of the test statistic
$T_{k}(\bm{X})\mid\{\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x}),\mathcal{Q}(\bm{X})=\mathcal{Q}(\bm{x})\}$
@@
is a truncated $\chi$-distribution $\mathrm{TC}(\mathrm{tr}(P_{k}),\mathcal{Z})$ with the degrees of freedom $\mathrm{tr}(P_{k})$ and the truncation region $\mathcal{Z}$. The truncation region $\mathcal{Z}$ is defined as
$\mathcal{Z}=\left\{z\in\mathbb{R}\mid\mathcal{A}({\bm{a}+\bm{b}z})=\mathcal{A}(\bm{x})\right\},\,\bm{a}=\mathcal{U}(\bm{x}),\,\bm{b}=\mathcal{V}(\bm{x}).$
The proof of Theorem 1 is deferred to Appendix B.1. By using the sampling distribution of the test statistic $T_{k}(\bm{X})$ conditional on $\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x})$ and $\mathcal{Q}(\bm{X})=\mathcal{Q}(\bm{x})$ in Theorem 1, we can define the selective $p$-value as
$p_{k}^{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0,k}}\left(T_{k}(\bm{X})\geq T_{k}(\bm{x})\mid\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x}),\mathcal{Q}(\bm{X})=\mathcal{Q}(\bm{x})\right).$ (17)
Theorem 2.
The selective $p$-value defined in (17) satisfies the property:
@@
$\mathbb{P}_{\mathrm{H}_{0,k}}\left(p_{k}^{\mathrm{selective}}\leq\alpha\mid\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x})\right)=\alpha,\ \forall\alpha\in(0,1).$
Then, the selective $p$-value also satisfies the following property of a valid $p$-value:
$\mathbb{P}_{\mathrm{H}_{0,k}}(p_{k}^{\mathrm{selective}}\leq\alpha)=\alpha,\ \forall\alpha\in(0,1).$
The proof of Theorem 2 is deferred to Appendix B.2. This theorem guarantees that the selective $p$-value is uniformly distributed under the null hypothesis $\mathrm{H}_{0,k}$, and thus can be used to conduct the valid statistical inference in (8) and (9). Figure 4 schematically illustrates the SI framework presented in this section.
@@
Once the truncation region $\mathcal{Z}$ is identified, the selective $p$-value in (17) can be easily computed by Theorem 1. Thus, the remaining task is reduced to identifying the truncation region $\mathcal{Z}$. In this study, to identify the truncation region $\mathcal{Z}$, we adopt the method based on parametric-programming technique proposed by Duy & Takeuchi (2022). For further details, refer to Appendix C.
Figure 4: Schematic illustration of the SI framework. A point in the data space $\mathbb{R}^{N}$ corresponds to a sequence with length $N$. The darkly shaded regions in the data space indicate that, if we input a point in these regions into the algorithm $\mathcal{A}$, the CP candidates are the same as $\mathcal{A}(\bm{x})$ obtained from the observed sequence $\bm{x}$. By conditioning on these regions and $\mathcal{Q}(\bm{X})=\mathcal{Q}(\bm{x})$, the conditional sampling distribution of the test statistic $T_{k}(\bm{X})$ is represented as a truncated $\chi$-distribution. Thus, selective $p$-values are defined based on the tail probability of such a truncated $\chi$-distribution.
@@ 5.2 Synthetic Data Experiments
Experimental setup.
In all synthetic experiments, we set window size $M\in\{512,1024\}$, the number of frequencies $D=\left\lfloor\frac{M}{2}\right\rfloor+1$, the length of sequence $N=M\cdot T$, where $T$ was specified for each experiment, the sampling rate $f_{s}=20480$, and each element of mean vector $\bm{s}$ as
@@
where frequencies $d_{1},d_{2},d_{3}\in\{0,...,D-1\}$ were randomly selected without replacement for each simulation, $A_{n}^{(d)}$ was defined for each experiment, and $\omega^{(d)}\in\left\{2\pi(\frac{f_{s}}{M})d\,\big{|}\,d=0,\dots,\left\lfloor\frac{M}{2}\right\rfloor\right\}$. We used BIC for the choice of penalty parameters $\bm{\beta}$ and $\gamma$ as indicated in Appendix A, and set the parameters of simulated annealing as $c_{0}^{+}=1000,\lambda^{+}=1.5,\eta(c_{0})=0.5$ and $\lambda=0.8$ in Section 3.4. After detecting CP candidates, a CP candidate location $\tau_{k}^{\text{det}}$ randomly selected from $\bm{\tau}^{\text{det}}$ was tested at the significance level $\alpha=0.05$.
In the experiments conducted to evaluate the type I error rate control, we generated $1000$ null sequences, which did not contain true CPs in the frequency domain, $\bm{x}=(x_{1},\dots,x_{N})^{\top}\sim\mathcal{N}\left(\bm{s},\sigma^{2}I_{N}\right)$, where $A_{n}^{(d)}=A^{(d)}$ was randomly sampled from $[0,1)$ for $d$ in each simulation, and $\sigma=1$, for each $T\in\{40,60,80,100\}$.
@@
$A_{n}^{(d)}=\begin{cases}A^{(d)}&\text{if $1\leq t\leq M\cdot t_{1}^{(d)}$},\\ \vspace{-3mm}\\ A^{(d)}+\Delta&\text{if $M\cdot t_{1}^{(d)}+1\leq t\leq M\cdot t_{2}^{(d)}$},\\ \vspace{-3mm}\\ A^{(d)}+2\Delta&\text{if $M\cdot t_{2}^{(d)}+1\leq t\leq T$},\end{cases}$
with $A^{(d_{1})},A^{(d_{2})},A^{(d_{3})}\in[0,1)$ which were randomly sampled in each simulation, $\left(t_{1}^{(d_{1})},t_{1}^{(d_{2})},t_{1}^{(d_{3})}\right)=(18,20,22),\left(t_{2}^{(d_{1})},t_{2}^{(d_{2})},t_{2}^{(d_{3})}\right)=(38,40,42)$, an intensity of the change $\Delta\in\{0.04,0.08,0.12,0.16\}$ and $\sigma=1$, for $T=60$. In each case, we ran $1000$ trials. Since we tested only when a CP candidate location was correctly detected, the power was defined as follows:
@@
We considered the CP candidate location $\tau_{k}$ to be correctly detected if it satisfied the following two conditions:
  • The set $\mathcal{D}_{k}$ of frequencies containing at least one CP candidate was a subset of $\{d_{1},d_{2},d_{3}\}$.
  • For $\mathcal{D}_{k}$ satisfying the above condition, either $\underset{d\in\mathcal{D}_{k}}{\min}{t_{1}^{(d)}}\leq\tau_{k}\leq\underset{d\in\mathcal{D}_{k}}{\max}{t_{1}^{(d)}}$ or $\underset{d\in\mathcal{D}_{k}}{\min}{t_{2}^{(d)}}\leq\tau_{k}\leq\underset{d\in\mathcal{D}_{k}}{\max}{t_{2}^{(d)}}$ held, that is, $\tau_{k}$ was detected within the true CP locations for the frequencies in $\mathcal{D}_{k}$.
Experimental results.
The results of experiments regarding the control of the type I error rate are shown in Figure 5. The Proposed, OC, and Bonferroni successfully controlled the type I error rate below the significance level, whereas the OptSeg-SI, OptSeg-SI-oc, and Naive could not. That was because the OptSeg-SI and OptSeg-SI-oc used $p$-values conditioned only on the dynamic programming algorithm, excluding the conditioning on simulated annealing, and the Naive employed conventional $p$-values without conditioning. Since the OptSeg-SI, OptSeg-SI-oc, and Naive failed to control the type I error rate, we omitted the analysis of their power. The results of power experiments are shown in Figure 6. Based on these results, the Proposed was the most powerful of all methods that controlled the type I error rate. The power of the OC was lower than that of the Proposed due to redundant conditions (see Appendix C for details). Furthermore, the Bonferroni method had the lowest power because it was a highly conservative approach that accounted for the huge number of all possible hypotheses. Additionally, we provide the computational time of the Proposed in both experiments and the information on the computer resources in Appendix D.2.
(a) $M=512$.
(b) $M=1024$.
Figure 5: Type I error rate.
(a) $M=512$.
(b) $M=1024$.
Figure 6: Power.
Additional experiments.
We also conducted the following experiments to investigate the robustness of the Proposed in terms of type I error rate control.
@@
In addition, a sensitivity study for the penalty parameter $\gamma$ within the objective function in (5) was also performed. These details and results are shown in Appendix D.3 and D.4, respectively.
@@ 5.3 Real Data Experiments
To demonstrate the practical applicability of the Proposed, we applied the Proposed, OC, and Naive to a real-world dataset. We used the set No.2 of the IMS bearing dataset, which is provided by the Center for Intelligent Maintenance Systems (IMS), University of Cincinnati (Qiu et al., 2006) and is available from the Prognostic data repository of NASA (Lee et al., 2007). The experimental apparatus consisted of four identical bearings (bearings 1, 2, 3, and 4) installed on a common shaft, driven at a constant rotation speed by an AC motor under applied radial loading. In this dataset, the vibration signals were measured using accelerometers until the outer race of bearing 1 failed at the end of the experiment, as shown in Figure 7. The signal with the sampling rate $f_{s}=20480$ was recorded for one second at intervals of 10 minutes over a period of about 7 days for each bearing. The analysis for the signal of bearing 1 in the frequency domain, as conducted by Gousseau et al. (2016), had revealed that enhancements of the spectral intensity were detected in harmonics of the characteristic frequency ($236$ Hz) associated with the outer race failure (Ball Pass Frequency Outer race, BPFO) on 3–4 days. Based on this previous study, we conducted CP candidate selection for the sensor data of bearing 1 in the frequency domain for two periods: 0.25–2.25 days when no significant changes in the spectra existed and 2.25–4.25 days when the BPFO harmonics exhibited the spectral amplification. For each period, we computed the DFT of $M=1024$ consecutive points in the 20480 samples and repeated the procedure $T=\text{2 days}/\text{10 min}=288$ times. As a result of the DFT, we obtained components for 513 frequencies (0, 20, …, 10240 Hz) and conducted CP candidate selection using 131 frequencies within the range of 1400–4000 Hz, which included the 6th to 16th harmonics of the characteristic frequency. For other parameters related to the CP candidate selection algorithm, please refer to the experimental setup in Section 5.2. Subsequently, we tested the detected CP candidate locations to evaluate whether each of them was a genuine CP location. The variance $\sigma^{2}$ for testing was estimated from the data on 0–0.25 days that did not contain significant spectral changes and was not used in any of the experiments. The results for the signal of bearing 1 on 0.25–2.25 days and 2.25–4.25 days are shown in Figure 8. In panel (a), the time variation of a frequency spectrum (1920 Hz) where a CP candidate location was falsely detected is shown for the period of 0.25–2.25 days. It shows that $p$-values of the Proposed and OC are above the significance level $0.05$, and therefore the result provides the validity of the inference, while $p$-value of the Naive is too small. In panel (b), the time variations of the 8th and 15th harmonics of the BPFO where CP candidate locations were correctly detected are presented for the period of 2.25–4.25 days. In this case, $p$-values of the Proposed are below the significance level $\frac{0.05}{2}=0.025$ decided by Bonferroni correction, thus it indicates that the inference is valid. In contrast, $p$-values of the OC are too large, due to the loss of power caused by the redundant conditions. In addition, since even the time sequences of the healthy bearings 2, 3, and 4 had been reported to indicate spectral amplification associated with the outer race fault in bearing 1 (Gousseau et al., 2016), we performed the same analysis for the three signals. The results are shown in Appendix D.5.
(a) Bearing 1.
(b) Bearing 2.
(c) Bearing 3.
(d) Bearing 4.
Figure 7: Vibration signals of the four bearings in the set No.2 of the IMS bearing dataset. Each signal was measured using an accelerometer for about 7 days until the outer race of bearing 1 failed. We used not only the signal of the damaged bearing 1 but also those of the other healthy bearings for the analysis because the spectral amplification associated with the failure of bearing 1 had been reported to be observed in the signals of all four bearings (Gousseau et al., 2016).
@@
(a) Inference on a falsely detected CP candidate location for 1920 Hz (around the 8th harmonic) on 0.25–2.25 days.
@@
(b) Inferences on truely detected CP candidate locations for 1880 Hz (the 8th harmonic, above) and 3540 Hz (the 15th harmonic, below) on 2.25–4.25 days.
@@
6 Conclusion, Limitation, and Future Work
In this paper, we developed a statistical inference method to quantify the statistical significance of detected CP locations in the frequency domain. Our proposed framework contributes to various fields where time-frequency analysis is widely employed, such as condition monitoring of machine systems using vibrational, electrical, and acoustic signals, and medical diagnosis based on biosignals. We conducted comprehensive experiments on both synthetic and real-world datasets. The results theoretically confirmed that our method provided an unbiased evaluation based on SI framework and demonstrated its superior performance compared to existing methods. However, an important limitation is the assumption that the noise follows an uncorrelated normal distribution, which is technically essential for deriving the conditional sampling distribution of the test statistic as a truncated $\chi$-distribution. In future work, we will extend our method to the case of multi-dimensional sequences. For instance, analyzing different types of time series obtained from multiple sensors would reveal significant changes in the system unattainable through the univariate approach, and reliability guarantee for the detections also provide a valuable future contribution.